OpenStudy (anonymous):

One canned juice drink is 25% orange juice; another is 10% orange juice . How many liters of each should be mixed together in order to get 15L that is 24% orange juice?

OpenStudy (anonymous):

for the first juice drink xLT for the second juice drink yLT. 25*x/100 +10*y/100= 24(x+y) x=14y and x+y =15LT 15y=15 y=1LT x=14LT

OpenStudy (anonymous):

We get (25x + 10y)/(x+y) = 24 and x+y = 15. Therefore, we plug in x+y = 15 and x = 15-y into our equation to get: 15*24 = 25 (15-y) +10y which equals 360 = 375 - 15y, or y = 1. If y = 1, then x= 14 The answer is 14L of 25% juice and 1L of 10% juice